How to find no. Of irrational rootts of


For eg. (x-1)(x-2)(3x-2)(3xplus1)=0

Grade:12

2 Answers

jitender lakhanpal
10 years ago

hi shubhum

the roots of above equation are

x = 1 x = 2 x = 2/3    x = -1/3

these are the only roots of this quartic equation and all of them are rational

there can at max be 4 roots of quartic equations

so no irrational roots

We here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

jitender

askiitians expert

jitender lakhanpal
10 years ago

pls provide your point of view or pls tell which part of the solution is wrong